MF Corp. has an ROE of 18% and a plowback ratio of 40%. The market capitalization rate is 13%. a. If the coming years earnings are expected to be $2.70 per share, at what price will the stock sell? (Do not round intermediate calculations. Round your answer to 2 decimal places.) Price b. What price do you expect MF shares to sell for in three years? (Do not round intermediate calculations. Round your answer to 2 decimal places.) Price
Answer #1

a.

growth rate or g=ROE* plowback ratio

=18%*40%

=7.20%

D1 or dividend in year 1=2.70*(1-40%)=1.62

price sell today=D1/(r-g)

=1.62/(13%-7.20%)

=27.93

b.

price in 3 years=27.93*(1+7.20%)^3=34.41

the above is answer..